OpenStudy (anonymous):

solve by writting the percent as a fraction fifth grade math q is 60% of 40 please put it in step form thanks

OpenStudy (anonymous):

I think you need to be clearer about what the question is asking. Can you rephrase, please?

OpenStudy (anonymous):

hold please

OpenStudy (anonymous):

i have to write the percent a a fraction the q is 60% of 40 is blank?

OpenStudy (anonymous):

would it be 3over 5?

OpenStudy (anonymous):

I honestly don't understand the question. However, I think what you might be asking is: By writing the percentage as a fraction, find out what 60% of 40 is. *Assuming* that's the question, you know that $$60\% = \frac{60}{100}$$ Therefore, 60% of 40 is: $$40\times\frac{60}{100} = \frac{40\times 60}{100}$$ The rest I leave up to you. :-)

OpenStudy (anonymous):

thanks for your help take care.